OpenStudy (anonymous):

Find the median for the following frequency distribution. x 20 21 22 23 24 25 26 27 28 29 30 f 4 2 1 3 6 2 5 8 4 1 3

OpenStudy (anonymous):

sorry the numbers are placed really weird.

OpenStudy (anonymous):

x 20 21 22 23 24 25 26 27 28 29 30 f 4 2 1 3 6 2 5 8 4 1 3

OpenStudy (anonymous):

I fixed the numbering. hope that helps.

OpenStudy (raden):

what's the total of freq... ?

OpenStudy (anonymous):

39

OpenStudy (campbell_st):

find the 20th score... that will give you the median score

OpenStudy (anonymous):

how do i find the 20th score?

OpenStudy (campbell_st):

the rule for the median is that if there are n scores then the median score is (n + 1)/2 score. so (39 + 1)/2 = 20th score

OpenStudy (raden):

when u calcul the sum of f1 + f2 + f3 + ... + f... u will found of the 20th score, is 4+2+1+3+6+2+5 so, the 20th score when the freq (f) = 5

OpenStudy (raden):

what's data for f=5 ?

OpenStudy (anonymous):

These are my options. A. 25 B. 26 C. 28 D. 27

OpenStudy (raden):

yea, i have looked ur answer smile to me :)

OpenStudy (anonymous):

haha what RadEn?

OpenStudy (campbell_st):

so the frequency of the scores 20 to 25 is 18 the frequency of the scores 20 to 26 is 23... so where do you think the 20th score would be?

OpenStudy (anonymous):

26?

OpenStudy (raden):

yes.. right

OpenStudy (anonymous):

ooh ok. but why does it have to add up to at least 20?

OpenStudy (raden):

to find location of the 20th data

OpenStudy (anonymous):

so what if it was a similar problem but instead of 20-30, it was 50-60. Would i try to find the 50th score?

OpenStudy (raden):

is that for group's datas

OpenStudy (anonymous):

x 50 51 52 53 54 55 56 57 58 59 60 f 14 8 7 5 8 12 9 7 10 16 14

OpenStudy (raden):

yeah, do same idea like before first, calcul the total of freq...

OpenStudy (anonymous):

110

OpenStudy (raden):

correct.. so, position of median will be at between of 55th data and 56th, right ?

OpenStudy (anonymous):

is it 55?

OpenStudy (anonymous):

if i add up to the 55th data that gives me 63

OpenStudy (anonymous):

sorry it's 56?

OpenStudy (anonymous):

the answer?

OpenStudy (raden):

yes, (56+56)/2 = 56

OpenStudy (anonymous):

I just did 14+8+7+5+8+12+9 and that gave me 63 but since that's up to 55. I just thought it was correct. haha

OpenStudy (raden):

lol, nopes... can more but dont less :P

OpenStudy (anonymous):

oh ok ok. so did i do it right? Also why did you add (55+56)/2 I don't understand why you added 56 to 55.

OpenStudy (raden):

56+56)/2 not 55+56)/2 because the sum of datas, is even number (110) so, no data is at the middle position right but between 2 numbers are the 55th data and 56th data

OpenStudy (raden):

for simple example ; there are numbers 1,2,3,4 so the position at median

OpenStudy (anonymous):

would that be 2+3 =5/2?

OpenStudy (raden):

at between 2 and 3, right ... 5/2 or 2.5

OpenStudy (anonymous):

oh i see. but the problems i had an odd amount of numbers, not even. which by the way I got right on my homework :) I didn't have a chance to say thank you

OpenStudy (anonymous):

sorry. i meant to say **the problems I had on my homework had an odd amount of numbers

OpenStudy (raden):

exept, if given of datas : 1,2,3,4,5 the value of median is 3 (because there are 5 datas (odd)) yes, good conclusion from u :)

OpenStudy (anonymous):

in the problems i had did the order of the frequencies matter?

OpenStudy (anonymous):

like did i have to arrange them from least to greatest?

OpenStudy (raden):

yes, alwasy order increasing of datas

OpenStudy (anonymous):

x 50 51 52 53 54 55 56 57 58 59 60 f 14 8 7 5 8 12 9 7 10 16 14 f: 5 7 7 8 8 9 10 12 14 14 16 (in order)

OpenStudy (raden):

do same idea... just difference of freq only...

OpenStudy (anonymous):

okay. thank you so much!

OpenStudy (raden):

very welcome, my friend ^^